# KVLink: Accelerating Large Language Models via Efficient KV Cache Reuse

### v3

@@ 1 Introduction
Large language models have demonstrated impressive capabilities across a broad array of applications—many of which involve processing contexts naturally divided into multiple segments. For example, in retrieval-augmented generation (RAG) [1, 2, 3], each retrieved document forms a distinct context chunk, while in multi-agent conversation scenarios [4, 5], outputs from different agents serve as separate segments. However, conventional architectures require LLMs to encode the entire concatenated context as a single unit before generating a response. This approach incurs high prefilling costs for long contexts and prevents the model from separately encoding and reusing precomputed representations (i.e., key-value states) for each segment. Consider RAG: for every query, the LLM encodes a large collection of retrieved documents. When different queries share common documents, the model redundantly re-encodes these identical texts, even though their content remains unchanged.
This inefficiency motivates us to explore an alternative strategy. Instead of re-encoding the entire concatenated context for every query, we propose precomputing the key-value (KV) states for each document or text segment independently, then reusing these precomputed states during inference. By encoding each segment (e.g., each retrieved document) separately and concatenating their KV states as needed, we can eliminate redundant computations and significantly improve efficiency. However, naively encoding each document independently and concatenating their KV caches during inference can lead to performance degradation due to train-test discrepancies. Prior work [6, 7, 8] has reported up to a 35% relative decrease in accuracy on QA tasks, which results from the discrepancy of position embeddings and missing cross attention between retrieved documents. To overcome this challenge, we introduce KVLink, an approach designed to bridge the gap between separately encoded segments and restore self-attention across documents. KVLink introduces two key enhancements: ❶ KV cache positional re-encoding. We adjust positional embeddings during inference to ensure that the stored KV caches align correctly with the positions required for a given query. ❷ Trainable cross-segment special tokens. To effectively “link” independently encoded segments, KVLink appends a small set of trainable tokens between each segment’s precomputed KV states before concatenation. The KV representations for these tokens are computed during inference. While the documents are independently encoded into KV cache, the link tokens can attend to all the preceding tokens, which helps restore self-attention across segments while introducing only minimal computational overhead.
@@
We validate the effectiveness and efficiency of our method through comprehensive experiments across diverse question-answering and text summarization datasets. We evaluate three backbone LLMs with different model scales, including Llama-3.2-1B, Llama-3.2-3B, and Llama-3.1-8B, showing that our approach consistently outperforms state-of-the-art baselines [6, 7, 9]. For example, KVLink surpasses the best baseline by 6.6% on Natural Question [10] and 7.3% on HotpotQA [11]. Even compared to the upper-bound model (i.e., LLM using conventional full encoding for inference), our method sacrifices only minimal performance while reducing the time-to-first-token latency by up to 96% by efficiently reusing precomputed KV caches.
Additionally, we investigate the integration of KVLink with KV cache compression techniques, as storing precomputed KV caches can incur substantial storage overhead. By incorporating methods such as LLMLingua [12] and AnLLMs [13], KVLink consistently outperforms baseline approaches. These results demonstrate its strong potential for practical, large-scale deployment.
@@ 2.1 Problem Formulation
However, the above approach often yields noticeable performance degradation, as LLMs are typically trained on fully concatenated sequences and each token attends to the preceding context. When we instead encode each document in a context-free manner, the model loses cross-document dependencies. Previous work also empirically demonstrates up to a 35% relative decrease in accuracy in question-answering tasks when each retrieved document is encoded into KV cache separately [7]. Our method aims to enable the LLM to produce high-quality outputs when the KV cache of each document is pre-computed independently by introducing two key components, ❶ KV cache positional re-encoding and ❷ cross-document reconnection with link tokens.
@@ 2.2 KV Cache Positional Re-encoding
[…]
@@
More specifically, we denote the hidden state of a token at position $i$ at a particular transformer layer as $\bm{x}_{i}\in\mathbb{R}^{d}$, where $d$ is the hidden dimension. The key vector is computed as $R_{i}W_{k}\bm{x}_{i}$ and the value vector is computed as $W_{v}\bm{x}_{i}$. Here $W_{\{k,v\}}\in\mathbb{R}^{d\times d}$ represents the weight matrices that project the hidden states into the key and value spaces, and $R_{i}\in[-1,1]^{d\times d}$ is the position-dependent rotation matrix used in RoPE. Because RoPE directly encodes positional information via these rotations, the KV cache can be stored without positional embeddings, i.e., $W_{\{k,v\}}\bm{x}_{i}$. At inference time, the key-value states of all documents are concatenated, and we apply the global rotary embedding for the KV states of each token appropriate to its location in the full sequence. This operation only introduces negligible time, ensuring the efficiency of our method. This is also consistent with previous methods [6, 8], which adjust the position encoding of separately encoded KV cache during inference.
@@ 2.4 Compressed KV Cache Linking
Although reusing pre-computed KV caches can greatly accelerate inference, it may introduce considerable overhead in storing these caches in a database and loading them onto the GPU. This overhead presents a major challenge for integrating cache reuse into real-world RAG deployments. For example, a 1,000-token document stored as UTF-8 text occupies only about 5KB, whereas the corresponding Llama3-8B KV cache requires roughly 131MB. To address this issue, we further explore combining our method with KV cache compression techniques. Specifically, we consider the following choices: ❶ LLMLingua [12], which compresses documents by token dropping; and ❷ AnLLMs [13], which fine-tunes the LLM to compress each sentence in the document into an anchor token with special attention masks.
@@ 3.1 Experiment Setup
We evaluate our method against three existing approaches: ❶ PromptCache [9], which directly reuses the precomputed KV cache of each retrieved document or context segment for model inference. Since the original PromptCache does not handle positional encoding mismatches, we enhance it by applying our positional re-encoding method (Section2.2) to ensure the stored KV cache aligns correctly with the concatenated input. ❷ CacheBlend [6], which concatenates the precomputed KV caches of retrieved documents and then recomputes the KV cache for a small number of selected tokens within the retrieved documents. Following the original implementation, we set the recomputation ratio to 18%. ❸ BlockAttention [7], which explicitly trains the model to handle separately encoded KV caches by fine-tuning on QA tasks where retrieved documents are encoded independently. For all baselines and our method, we use Llama-3.2-1B-Instruct and Llama-3.2-3B-Instruct [15] as the backbone models. Since BlockAttention and our method require fine-tuning, we ensure fair comparison by using the same data mixture and training hyperparameters for both. Further implementation details for each baseline are provided in Appendix A.3.
@@
We adopt Llama-3.2-1B-Instruct, Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct as the backbone models, fine-tuning them for 6,000 steps using a global batch size of 64 across 8$\times$×H100 GPUs. We construct the training dataset by mixing the training sets of 2WikiMQA , TriviaQA , pretraining data from FineWeb [16], and Tülu 3 [17]. Further details on data preprocessing, dataset mixture, and training configurations are provided in Appendix A.1 and  A.2. For our method, we train three versions, each appending 0, 1, or 5 link tokens to every document or context segment, denoted as KVLink0, KVLink1, and KVLink5, respectively.
@@
To evaluate the model performance when using separately encoded documents or context segments, we focus on retrieval-augmented question answering tasks including NaturalQuestions [10], 2WikiMQA [18], TriviaQA [19], HotpotQA [11], and MuSiQue [20]. For NaturalQuestions, we adopt the evaluation protocol from Liu et al. [21], where the LLM is given a question along with a set of 10 retrieved documents. The document that contains the correct answer is systematically placed at each of the 10 possible positions across separate evaluation runs. The final accuracy is reported as the average over these 10 evaluations. For 2WikiMQA, HotpotQA, and MuSiQue, we utilize the originally provided retrieved documents for evaluation. For TriviaQA we retrieve 10 documents using Contriever [22] following the setting in BlockAttention. In all cases, documents are encoded separately into KV cache.
Additionally, following CacheBlend, we evaluate on two text summarization datasets including MultiNews [23] and Samsum [24]. The LLM is given several in-context examples per instance, each separately encoded into KV cache. We report the RougeL score [25] as the metric. Notably, our evaluation datasets cover all those used by our baselines, BlockAttention and CacheBlend, ensuring a comprehensive and fair comparison.
@@
To ensure our method does not degrade the model’s original capabilities, we evaluate it on a range of reasoning and instruction-following benchmarks, including IFEval [26], GSM8K [27], MMLU [28], ARC-Challenge [29], ARC-Easy [29], PiQA [30], SciQ [31], Winogrande [32], and HellaSwag [33]. We report the accuracy on each dataset. For IFEval, we report both strict prompt-level accuracy (IFEval-P) and instruction-level accuracy (IFEval-I). More details of the evaluation configuration are available in Appendix A.2.
Finally, to further reduce cache-storage and loading overhead, and thereby make cache reuse practical, we evaluate reuse with compressed KV caches. We report QA accuracy under various compression methods and assess the effectiveness of KVLink when operating on the compressed caches. The only difference from the first experiment is that we use compressed, rather than original, KV cache.
@@ 3.3 Inference Efficiency Evaluation
A key objective of KVLink is to reduce the computational overhead incurred by repeatedly encoding long contexts. To demonstrate these efficiency gains, we measure the Time-to-First-Token (TTFT) when reusing the precomputed KV cache for ten documents of varying lengths. For a realistic assessment, we include the overhead of loading the precomputed KV caches onto the GPU. Each measurement is averaged over 100 runs, with an initial 10 warm-up trials to eliminate memory allocation overhead, following the setup in  [34].
@@ 4 Related Work
Efficient inference for LLMs. As model sizes grow, serving LLMs efficiently becomes increasingly challenging. Previous work has tackled this via model pruning [35, 36, 37, 38], quantization [39, 40, 41, 42, 43, 44], and optimized decoding algorithms such as non-autoregressive decoding [45], speculative decoding [46, 47, 48], or early-exiting [49, 50]. Like these approaches, our method also aims to enhance LLM inference efficiency by reducing inference-time computation.
@@
KV cache reuse. Recent work has explored reusing precomputed KV caches across multiple queries, often focusing on prefix-only reuse [51, 52, 53]. A few methods extend reuse to arbitrary positions, aiming for greater efficiency. PromptCache[9] allows KV cache reuse with discontinuous position IDs but ignores cross-chunk attention, causing performance drops. It also relies on Prompt Markup Language, limiting flexibility across tasks. CacheBlend reduces positional mismatches by selectively recomputing caches for tokens with the largest value-state variance; however, it suffers from performance degradation and the non-trivial cost of recomputing 10%–20% of tokens (plus all tokens at the first layer). BlockAttention[7] removes positional embeddings by re-encoding KV caches and directly fine-tunes on block-attention data, but lacks strategies to better connect reused contexts or refine data processing, leaving performance below ideal levels.
@@
KV compression. While our method focuses on accelerating inference by reusing KV caches, a complementary line of work aims to reduce memory overhead through cache eviction and quantization. For instance,  [54, 55, 56, 57] introduces an eviction policy that significantly lowers the memory footprint during generation, while  [58, 42, 39, 41, 59, 60, 61] investigate quantizing KV caches to minimize storage requirements with only minor performance loss. Notably, our approach can be seamlessly integrated with these state-of-the-art eviction and quantization strategies, thereby addressing both speed and memory efficiency.
@@
Retrieve-augmented generation. Retrieval-augmented generation (RAG) is widely used to enhance the generation capabilities of language models by retrieving supporting documents. Retrievers such as dense retrievers or BM25 [22, 62, 63] are typically employed to find the most relevant documents for a given user query or context. In addition to simply concatenating the retrieved text with the input [64], there are also other methods for integrating the retrieved knowledge into the models. For instance, LongMem [65] encodes documents into caches and uses a trained side net to incorporate these knowledge caches into the context for long-context tasks. With the advances of LLMs across various NLP tasks, certain studies have focused on improving RAG with LLM [66, 67]. Some research also explores how to enhance the robustness of LLMs when processing retrieved knowledge. For example, RAFT [68] provides a fine-tuning strategy that strengthens LLMs against noisy contexts in domain-specific RAG.

### v4

@@ 1 Introduction
Large language models have demonstrated impressive capabilities across a broad array of applications—many of which involve processing contexts naturally divided into multiple segments. For example, in retrieval-augmented generation (RAG) [gao2023retrieval, wang2024searching, li2024retrieval], each retrieved document forms a distinct context chunk, while in multi-agent conversation scenarios [wu2023autogen, liu2024dynamic], outputs from different agents serve as separate segments. However, conventional architectures require LLMs to encode the entire concatenated context as a single unit before generating a response. This approach incurs high prefilling costs for long contexts and prevents the model from separately encoding and reusing precomputed representations (i.e., key-value states) for each segment. Consider RAG: for every query, the LLM encodes a large collection of retrieved documents. When different queries share common documents, the model redundantly re-encodes these identical texts, even though their content remains unchanged.
This inefficiency motivates us to explore an alternative strategy. Instead of re-encoding the entire concatenated context for every query, we propose precomputing the key-value (KV) states for each document or text segment independently, then reusing these precomputed states during inference. By encoding each segment (e.g., each retrieved document) separately and concatenating their KV states as needed, we can eliminate redundant computations and significantly improve efficiency. However, naively encoding each document independently and concatenating their KV caches during inference can lead to performance degradation due to train-test discrepancies. Prior work [yao2024cacheblend, sun2024block, zhang2024attention] has reported up to a 35% relative decrease in accuracy on QA tasks, which results from the discrepancy of position embeddings and missing cross attention between retrieved documents. To overcome this challenge, we introduce KVLink, an approach designed to bridge the gap between separately encoded segments and restore self-attention across documents. KVLink introduces two key enhancements: ❶ KV cache positional re-encoding. We adjust positional embeddings during inference to ensure that the stored KV caches align correctly with the positions required for a given query. ❷ Trainable cross-segment special tokens. To effectively “link” independently encoded segments, KVLink appends a small set of trainable tokens between each segment’s precomputed KV states before concatenation. The KV representations for these tokens are computed during inference. While the documents are independently encoded into KV cache, the link tokens can attend to all the preceding tokens, which helps restore self-attention across segments while introducing only minimal computational overhead.
@@
We validate the effectiveness and efficiency of our method through comprehensive experiments across diverse question-answering and text summarization datasets. We evaluate three backbone LLMs with different model scales, including Llama-3.2-1B, Llama-3.2-3B, and Llama-3.1-8B, showing that our approach consistently outperforms state-of-the-art baselines [yao2024cacheblend, sun2024block, gim2024prompt]. For example, KVLink surpasses the best baseline by 6.6% on Natural Question [kwiatkowski2019natural] and 7.3% on HotpotQA [yang2018hotpotqa]. Even compared to the upper-bound model (i.e., LLM using conventional full encoding for inference), our method sacrifices only minimal performance while reducing the time-to-first-token latency by up to 96% by efficiently reusing precomputed KV caches.
Additionally, we investigate the integration of KVLink with KV cache compression techniques, as storing precomputed KV caches can incur substantial storage overhead. By incorporating methods such as LLMLingua [pan2024llmlingua] and AnLLMs [pang2024anchor], KVLink consistently outperforms baseline approaches. These results demonstrate its strong potential for practical, large-scale deployment.
@@ 2.1 Problem Formulation
However, the above approach often yields noticeable performance degradation, as LLMs are typically trained on fully concatenated sequences and each token attends to the preceding context. When we instead encode each document in a context-free manner, the model loses cross-document dependencies. Previous work also empirically demonstrates up to a 35% relative decrease in accuracy in question-answering tasks when each retrieved document is encoded into KV cache separately [sun2024block]. Our method aims to enable the LLM to produce high-quality outputs when the KV cache of each document is pre-computed independently by introducing two key components, ❶ KV cache positional re-encoding and ❷ cross-document reconnection with link tokens.
@@ 2.2 KV Cache Positional Re-encoding
The first issue with separately encoding documents is the position mismatch during inference. Modern LLMs typically use Rotary Positional Encoding (RoPE) [su2024roformer], where each token is assigned a distinct positional embedding based on its position in the full sequence. However, when documents are encoded independently, tokens are indexed within their own document, ignoring their actual position in the full concatenated input. Take Figure 1 as an example. Since we pre-compute the KV cache for each document separately, the second token in $\mathrm{Doc_{b}}$ is assigned position index $2$ when computing its KV cache. However, when we concatenate $\mathrm{Doc_{b}}$ with $\mathrm{Doc_{a}}$ during inference (e.g., when processing query $\bm{q}_{1}$), the actual position of that token should be $|\texttt{$\mathrm{Doc_{a}}$}|+2$ where $|\texttt{$\mathrm{Doc_{a}}$}|$ is the length of the first document. Since the KV cache of $\mathrm{Doc_{b}}$ was precomputed without awareness of this global position shift, the LLM applies incorrect positional embeddings, leading to erroneous attention computations.
@@
More specifically, we denote the hidden state of a token at position $i$ at a particular transformer layer as $\bm{x}_{i}\in\mathbb{R}^{d}$, where $d$ is the hidden dimension. The key vector is computed as $R_{i}W_{k}\bm{x}_{i}$ and the value vector is computed as $W_{v}\bm{x}_{i}$. Here $W_{\{k,v\}}\in\mathbb{R}^{d\times d}$ represents the weight matrices that project the hidden states into the key and value spaces, and $R_{i}\in[-1,1]^{d\times d}$ is the position-dependent rotation matrix used in RoPE. Because RoPE directly encodes positional information via these rotations, the KV cache can be stored without positional embeddings, i.e., $W_{\{k,v\}}\bm{x}_{i}$. At inference time, the key-value states of all documents are concatenated, and we apply the global rotary embedding for the KV states of each token appropriate to its location in the full sequence. This operation only introduces negligible time, ensuring the efficiency of our method. This is also consistent with previous methods [yao2024cacheblend, zhang2024attention], which adjust the position encoding of separately encoded KV cache during inference.
@@ 2.4 Compressed KV Cache Linking
Although reusing pre‑computed KV caches can greatly accelerate inference, it may introduce considerable overhead in storing these caches in a database and loading them onto the GPU. This overhead presents a major challenge for integrating cache reuse into real-world RAG deployments. For example, a 1,000‑token document stored as UTF‑8 text occupies only about 5KB, whereas the corresponding Llama3‑8B KV cache requires roughly 131MB. To address this issue, we further explore combining our method with KV cache compression techniques. Specifically, we consider the following choices: ❶ LLMLingua [pan2024llmlingua], which compresses documents by token dropping; and ❷ AnLLMs [pang2024anchor], which fine-tunes the LLM to compress each sentence in the document into an anchor token with special attention masks.
@@ 3.1 Experiment Setup
We evaluate our method against three existing approaches: ❶ PromptCache [gim2024prompt], which directly reuses the precomputed KV cache of each retrieved document or context segment for model inference. Since the original PromptCache does not handle positional encoding mismatches, we enhance it by applying our positional re-encoding method (Section2.2) to ensure the stored KV cache aligns correctly with the concatenated input. ❷ CacheBlend [yao2024cacheblend], which concatenates the precomputed KV caches of retrieved documents and then recomputes the KV cache for a small number of selected tokens within the retrieved documents. Following the original implementation, we set the recomputation ratio to 18%. ❸ BlockAttention [sun2024block], which explicitly trains the model to handle separately encoded KV caches by fine-tuning on QA tasks where retrieved documents are encoded independently. For all baselines and our method, we use Llama-3.2-1B-Instruct and Llama-3.2-3B-Instruct [meta2024llama] as the backbone models. Since BlockAttention and our method require fine-tuning, we ensure fair comparison by using the same data mixture and training hyperparameters for both. Further implementation details for each baseline are provided in Appendix A.3.
@@
[…]
@@
To evaluate the model performance when using separately encoded documents or context segments, we focus on retrieval-augmented question answering tasks including NaturalQuestions [kwiatkowski2019natural], 2WikiMQA [ho2020constructing], TriviaQA [joshi2017triviaqa], HotpotQA [yang2018hotpotqa], and MuSiQue [trivedi2022musique]. For NaturalQuestions, we adopt the evaluation protocol from liu2024lost, where the LLM is given a question along with a set of 10 retrieved documents. The document that contains the correct answer is systematically placed at each of the 10 possible positions across separate evaluation runs. The final accuracy is reported as the average over these 10 evaluations. For 2WikiMQA, HotpotQA, and MuSiQue, we utilize the originally provided retrieved documents for evaluation. For TriviaQA we retrieve 10 documents using Contriever [izacard2021unsupervised] following the setting in BlockAttention. In all cases, documents are encoded separately into KV cache.
Additionally, following CacheBlend, we evaluate on two text summarization datasets including MultiNews [fabbri2019multi] and Samsum [gliwa2019samsum]. The LLM is given several in-context examples per instance, each separately encoded into KV cache. We report the RougeL score [lin2004rouge] as the metric. Notably, our evaluation datasets cover all those used by our baselines, BlockAttention and CacheBlend, ensuring a comprehensive and fair comparison.
@@
To ensure our method does not degrade the model’s original capabilities, we evaluate it on a range of reasoning and instruction-following benchmarks, including IFEval [zhou2023instruction], GSM8K [cobbe2021training], MMLU [hendrycks2020measuring], ARC-Challenge [clark2018think], ARC-Easy [clark2018think], PiQA [bisk2020piqa], SciQ [welbl2017crowdsourcing], Winogrande [sakaguchi2021winogrande], and HellaSwag [zellers2019hellaswag]. We report the accuracy on each dataset. For IFEval, we report both strict prompt-level accuracy (IFEval-P) and instruction-level accuracy (IFEval-I). More details of the evaluation configuration are available in Appendix A.2.
Finally, to further reduce cache‑storage and loading overhead, and thereby make cache reuse practical, we evaluate reuse with compressed KV caches. We report QA accuracy under various compression methods and assess the effectiveness of KVLink when operating on the compressed caches. The only difference from the first experiment is that we use compressed, rather than original, KV cache.
@@ 3.3 Inference Efficiency Evaluation
A key objective of KVLink is to reduce the computational overhead incurred by repeatedly encoding long contexts. To demonstrate these efficiency gains, we measure the Time-to-First-Token (TTFT) when reusing the precomputed KV cache for ten documents of varying lengths. For a realistic assessment, we include the overhead of loading the precomputed KV caches onto the GPU. Each measurement is averaged over 100 runs, with an initial 10 warm-up trials to eliminate memory allocation overhead, following the setup in  [xiao2024duoattention].
@@ 4 Related Work
Efficient inference for LLMs. As model sizes grow, serving LLMs efficiently becomes increasingly challenging. Previous work has tackled this via model pruning [men2024shortgpt, sreenivas2024llm, xiasheared, houinstruction], quantization [lin2024awq, van2024gptvq, frantar2022gptq, xiao2023smoothquant, yao2022zeroquant, dettmers2022gpt3], and optimized decoding algorithms such as non-autoregressive decoding [santilli2023accelerating], speculative decoding [kim2024speculative, elhoushi2024layer, miao2023specinfer], or early-exiting [he2021magic, kong2022accelerating]. Like these approaches, our method also aims to enhance LLM inference efficiency by reducing inference-time computation.
@@
KV cache reuse. Recent work has explored reusing precomputed KV caches across multiple queries, often focusing on prefix-only reuse [jin2024ragcache, liu2024optimizing, zheng2023efficiently]. A few methods extend reuse to arbitrary positions, aiming for greater efficiency. PromptCache[gim2024prompt] allows KV cache reuse with discontinuous position IDs but ignores cross-chunk attention, causing performance drops. It also relies on Prompt Markup Language, limiting flexibility across tasks. CacheBlend reduces positional mismatches by selectively recomputing caches for tokens with the largest value-state variance; however, it suffers from performance degradation and the non-trivial cost of recomputing 10%–20% of tokens (plus all tokens at the first layer). BlockAttention[sun2024block] removes positional embeddings by re-encoding KV caches and directly fine-tunes on block-attention data, but lacks strategies to better connect reused contexts or refine data processing, leaving performance below ideal levels. TurboRAG [lu2024turborag] introduces two special tokens to mark the boundaries of reused caches. However, it still falls short in effectively reconnecting reused caches, similar to BlockAttention. A detailed discussion is provided in Appendix A.8.
@@
KV compression. While our method focuses on accelerating inference by reusing KV caches, a complementary line of work aims to reduce memory overhead through cache eviction and quantization. For instance,  [zhang2024h2o, oren2024transformers, xiao2023efficient, ge2023model] introduces an eviction policy that significantly lowers the memory footprint during generation, while  [liu2024kivi, xiao2023smoothquant, lin2024awq, frantar2022gptq, kim2023squeezellm, zhao2024atom, sheng2023flexgen] investigate quantizing KV caches to minimize storage requirements with only minor performance loss. Notably, our approach can be seamlessly integrated with these state-of-the-art eviction and quantization strategies, thereby addressing both speed and memory efficiency.
@@
Retrieve-augmented generation. Retrieval-augmented generation (RAG) is widely used to enhance the generation capabilities of language models by retrieving supporting documents. Retrievers such as dense retrievers or BM25 [izacard2021unsupervised, lee2020learning, robertson2009probabilistic] are typically employed to find the most relevant documents for a given user query or context. In addition to simply concatenating the retrieved text with the input [izacard2020leveraging], there are also other methods for integrating the retrieved knowledge into the models. For instance, LongMem [wang2023augmenting] encodes documents into caches and uses a trained side net to incorporate these knowledge caches into the context for long-context tasks. With the advances of LLMs across various NLP tasks, certain studies have focused on improving RAG with LLM [cuconasu2024power, asai2023self]. Some research also explores how to enhance the robustness of LLMs when processing retrieved knowledge. For example, RAFT [zhang2024raft] provides a fine-tuning strategy that strengthens LLMs against noisy contexts in domain-specific RAG.